OpenStudy (anonymous):

find h(x,y) = g(f(x,y)) if g(t) = t^2 + sqrt(t) and f(x,y) = 8x+7y-6

OpenStudy (anonymous):

So I don't know where to start. Please help!

OpenStudy (anonymous):

not really sure, but it looks like it should be \[\large g(f(x,y))=g(8x+7y-6)=(8x+7y-6)^2+\sqrt{8x+7y-6}\]

OpenStudy (anonymous):

thanks. I think I get it now.